Solve the system.
5x + 5y – 4z = 13
5x + 3y + z = –3
3x – 2y – z = –10

Answer by Alan3354(69443)

Solve the system.
5x + 5y – 4z = 13
5x + 3y + z = –3
3x – 2y – z = –10
-------------------
Add eqns 2 & 3
---
Multiply the 2nd eqn by 4 and add to the 1st eqn.
z is eliminated, leaving 2 eqns in x & y.
Can you do the rest?